OpenStudy (anonymous):

A souvenir maker wants to create a scale model of the Eiffel Tower. The Eiffel Tower is 324 meters tall and has a base with dimensions 125 meters by 125 meters. The model will rest on a base with dimensions 12.5 cm by 12.5 cm. How tall will the model be in centimeters?

OpenStudy (nurali):

125/12.5 = 10 so the scale is 10cm:1m 324/10 = 32.4cm

OpenStudy (anonymous):

0.21 cm 20.74 cm 0.324 cm 32.40 cm

OpenStudy (nurali):

Last one.

OpenStudy (anonymous):

thanks

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

4. A souvenir maker wants to create a scale model of the Empire State Building. The Empire State Building is 1,472 feet tall and has a base with dimensions 286 ft by 286 ft. If the model is 6 in. tall, what are the approximate dimensions of its base in inches? (Round your answer to the nearest tenth of an inch.) (1 point) 1.2 in. by 1.2 in .1 in. by 1 in. 0.2 in. by 0.2 in. 2.3 in. by 2.3 in.

OpenStudy (anonymous):

5. Is (3, 13) a solution of the equation y = 4x? (1 point)yes or no

OpenStudy (nurali):

1.2 by 1.2 No

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

9. –4 = g + 1 (1 point)g = –4g = 4g = –5g = –3

OpenStudy (anonymous):

)g = –4 g = 4 g = –5 g = –3

OpenStudy (anonymous):

oh ok